OpenStudy (anonymous):

Find, in simplest form, the coefficient of x^r in the expansion, in ascending powers of x, of [1/(1-x)^3]

OpenStudy (anonymous):

(1-x)^-3

OpenStudy (anonymous):

use binomial theorem (1 + (-x))^-3 = inf ∑(-3Ck) (-x)^n k=0 now it's just the matter of expanding (-3Ck)

OpenStudy (anonymous):

typo, should be inf ∑(-3Cn) (-x)^n n=0

OpenStudy (anonymous):

can you take it from here?

OpenStudy (anonymous):

sorry i dont understand the ∑(-3Cn) (-x)^n, im suppose to use the [(n)(n-1)..(n-r+1)]/r! formula

OpenStudy (anonymous):

yes, take's exactly what you use. (-3Cn) = (-3)(-4)(-5) ... (-3-n+1) / n! = (-3)(-4)(-5) ... (-2-n) / n! = (-1)^n (3*4*5 * ... * (n+2) ) / n! = (-1)^n (1*2*3*4*5 *... * n (n+1) (n+2) ) /(1*2* n!) = (-1)^n (1/2) (n+1) (n+2)

OpenStudy (anonymous):

huhm.. seems like i'm off by a constant 1/2. Hold on, let me check my algebra again

OpenStudy (anonymous):

yes im self-studying

OpenStudy (anonymous):

@HatcrewS are you trying to find the power expansion form?

OpenStudy (anonymous):

im trying to find the general term for the coefficient of the expansion, the answer is [(r+1)(r+2)]/2

OpenStudy (anonymous):

which is exactly what i got O.O

OpenStudy (anonymous):

hmm but you got the (-1)^n >_< and i still don't understand how you get (-1)^n, how does the ^n come about ? I just memorise that if the sequence is negative I can take out -1 and make it become (-1)^n

OpenStudy (anonymous):

you can eliminate the (-1)^n by having another (-1)^n from (-x), then somehow making it (-1)^2n, which is always positive i guess :/

OpenStudy (anonymous):

that is because if you multiply a negative odd number of times, you'll get a negative number. And if you multiply a negative number an even number of times, then it's positive. (-1)^n determines the sign of the coefficient

OpenStudy (anonymous):

ah yes, (-1)^n * (-1)^n at the end will always end up positive. I didn't simplify all the way through